Youve observed the following returns on Doyscher Corporations stock over the past five years: -24.3 percent, 13.2 percent,
Answer #1

Average return=Total return/Total time period

=(-24.3+13.2+29.4+2.1+21.1)/5

=8.3%

Return (Return-Average return)^2
-24.3 (-24.3-8.3)^2=1062.76
13.2 (13.2-8.3)^2=24.01
29.4 (29.4-8.3)^2=445.21
2.1 (2.1-8.3)^2=38.44
21.1 (21.1-8.3)^2=163.84
Total=1734.26%

Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[1734.26/(5-1)]^(1/2)

=[1734.26/4]^(1/2)

=20.82%(Approx).

Variance=Standard deviation^2

=0.043357(Approx).